1.a) Consider an agent who will live for two periods with utility function U(x1, x2) = x1 * x * . The agent receives incomesPlease help me with this question, thank you so much!
Answer #1

Given -

r=10 %, income received in period 1 is $10 and in period 2 is $10. Optimal condition will be fulfilled at the point where slope of intertemporal consumption curve is equal to slope of utility function, that is (1+r)=MUx1/ MUx2, where x1 and x2 are consumption in period 1 and period 2 respectively.

Ginen, ITTI I = $10 I2 = $10. 8 = 10% U(X, H₂) = 24 x 3/4 optimal condition - (1+r)= MUXI Mitry slope of mura | Intertempora

80, aut At the optimal level, G+ C = I + I2 Tomta dojot tr u ller, te rochastol l Hure, Mt 2 = 1 + I2 Consumption in tr 1st p

Diagram

Diagram yam de (consump inand Period Budget constraint x + 2 y = I + Iz ty tr Slope = (1+r) Ę Optimal Level - muay - (ltr) =